2018 data update (34/n)
[Orig: August 10, 2018]
Things look to be calming down over the past week, though still trending slightly above normal. A small change with OGG points as we removed a comparison star that was occasionally falling on a hot pixel.
Note from previous post that still applies: The gray vertical line indicates the time of the change in configuration. There are still some remaining issues that we need to address for data taken primarily during the first week of June (where you see a gap in data) and hopefully we will be able to update this soon with those measurements. Also, all data taken after this time have an unknown vertical offset. As such, these measurements are set to be consistent with the flux at the end of May until we can better assess what, if anything, had changed during the gap.
~Tabby and team
PS: These observations are happening because of the wonderful backers of our 2016 Kickstarter project. The Kickstarter campaign has ended, but we are still accepting donations to purchase additional observing time on the LCO 0.4m network. Thanks in advance for your support!